# Beyond the Leaderboard: Rethinking Medical Benchmarks for Large Language Models

### v1

@@ Appendix B Quantitative Results
This appendix provides a summary of the quantitative analysis conducted on the 53 medical LLM benchmarks.
B.1 Publication Year and Task Focus of Benchmarks
@@
Annual Trend. As shown in Figure 4, the publication years of the 53 benchmarks we collected. A clear turning point can be observed in 2024, when the volume of medical benchmark publications significantly increased. Prior to 2024, only 15 benchmarks were released. The brief surge in 2019, driven by the rise of deep learning and Transformer models, catalyzed short-term growth in medical AI and a peak in benchmark publications. However, no substantial upward trend was evident before 2024.
Figure 4: Publication Year Distribution of the 53 Medical Benchmarks
In contrast, with the emergence of large models such as ChatGPT and Med-PaLM in 2024, there was an explosive growth in medical NLP and multimodal benchmarks. A total of 26 medical benchmarks were published in 2024, surpassing the cumulative number of the previous six years. As of July 10, 2025, 12 new benchmarks were already released in the first half of 2025, showing a steady growth trend.
@@
In summary, as illustrated in Figure 6 nearly half (47%) of the existing medical benchmarks focus on assessing models’ text comprehension, NLP, and VQA abilities. This indicates that most medical benchmarks historically centered around structured or unstructured medical text understanding. However, in recent years, there has been an increasing trend towards medical question answering and clinical NLP, with the highest number of such benchmarks being published in the first half of 2025.
Figure 6: Task Type Distribution of the 53 Medical Benchmarks
B.2 Natural Language
Figure 7 presents the distribution of natural language usage across the 53 collected benchmarks. It is evident that English dominates, being used in 42 of the benchmarks (approximately 79%), establishing itself as the primary language for evaluating medical LLMs. Chinese ranks second, appearing in 17 benchmarks (approximately 32%), indicating a notable level of development in Chinese medical AI. In contrast, other languages such as Arabic, Russian, Portuguese, Japanese, and Hindi—despite being spoken by large global populations—remain underrepresented in publicly available benchmarks, with each language appearing in only 2 to 6 benchmarks on average. This highlights that multilingual medical AI remains in its early stages.
@@ B.2 Natural Language
B.3 Medical Diversity
Distribution of Diseases. Figure 8 illustrates the distribution of disease categories represented in the 53 medical benchmarks, categorized according to the first 23 chapters of the ICD-11 classification system (with Certain conditions originating in the perinatal period and Developmental anomalies excluded due to the absence of relevant benchmarks). The analysis reveals that current research on medical LLMs is heavily concentrated in disease areas such as Diseases of the musculoskeletal system or connective tissue (16 benchmarks), Diseases of the skin (15), Diseases of the respiratory system (14), and Neoplasms (13). These categories typically have abundant publicly available imaging data (e.g., X-rays, CT scans, pathology slides), which facilitates the development of standardized evaluation protocols based on both image and text modalities.
@@ B.3 Medical Diversity
B.4 Benchmark Performance Heatmap Example
@@
The heatmap in Figure 10 visualizes the performance of 14 out of 53 of the medical LLM benchmarks evaluated evaluated against the five-phase MedCheck framework. Each row represents a benchmark, and columns correspond to the framework’s phases. The color intensity reflects the score rate (%), highlighting the comparative strengths and weaknesses of each benchmark and revealing systemic trends.
B.5 Statistics About Phase I: Design and Conceptualization
@@
As shown in Figure 11, the average normalized score across benchmarks is 74.3% (17.8 out of 24), with a mean of 1.5 points per criterion, indicating general compliance with baseline expectations. However, performance on several key items remains suboptimal and warrants further attention.
Figure 11: Scoring Performance of the 53 Benchmarks in the Design and Conceptualization Stage
Clearly specified application scenarios. To ensure practical relevance, benchmarks should clearly define their intended clinical or biomedical research scenarios. Such grounding facilitates real-world applicability and aligns evaluation with deployment needs. As illustrated in Figure 12, 66% of benchmarks specify concrete application scenarios, articulating clinical value and expected utility. In contrast, 19% provide only general references to decision-support contexts, and 15% fail to specify any scenario, reducing interpretability and downstream usability.
@@
Involvement of domain experts. Expert input is a critical component in medical AI development. It ensures the appropriateness of task design, data annotation quality, and answer validity. However, as shown in Figure 13, only 25% of benchmarks explicitly report domain expert involvement, often without detailing the roles or qualifications. 17% do not mention expert participation, raising concerns about the benchmarks’ clinical validity. Future benchmark design should adopt standardized expert reporting practices to enhance credibility, transparency, and methodological rigor.
@@
Alignment with medical standards. Adherence to international medical standards (e.g., ICD-11, SNOMED CT, LOINC) is essential for achieving semantic interoperability, annotation consistency, and cross-system comparability. As Figure 14 indicates, 53% of benchmarks do not clarify whether their data conforms to such standards. This omission undermines the benchmark’s reusability and weakens the comparability of evaluation outcomes. Without standard-based structure and terminology, label ambiguity and inconsistent categorization may arise, which compromises the reliability of model performance evaluation.
@@
Multidimensional evaluation design. In medical AI, a model’s performance cannot be solely measured by accuracy. Outputs that are correct but unsafe, incomplete, or biased may have detrimental consequences. Multidimensional evaluation frameworks are required to assess whether models’ output is not only correct but also reliable. As presented in Figure 15, 36% of benchmarks use only a single metric, typically accuracy, and 21% mention other metrics without detailed implementation. Metrics such as uncertainty estimation, answer refusal, and harmful output detection are absent from most benchmarks, despite their relevance in high-stakes medical decision-making.
@@
Consideration of safety and fairness. Robust evaluation frameworks must account for safety and fairness, especially in sensitive domains like healthcare. As shown in Figure 16, only 15% of benchmarks include explicit safety or fairness evaluations, while 38% acknowledge their importance without implementation. The remaining 47% omit these aspects entirely. Neglecting safety assessments increases the risk of misleading or overconfident outputs, which may harm clinical decisions. The lack of fairness evaluation may propagate systematic bias across gender, age, race, or disease distribution, exacerbating healthcare disparities. To build trustworthy and ethical medical AI, safety and fairness must be treated as core evaluation dimensions.
@@
B.6 Statistics About Phase II: Dataset Construction and Management
This phase constitutes the empirical foundation of the benchmark, with an emphasis on constructing datasets that are realistic, diverse, representative, and ethically sourced. Key elements include stringent procedures for data quality assurance, privacy protection, and contamination prevention. The dataset sources and quality were assessed across five dimensions: transparency and traceability of data provenance, reliability of data sources, and data authenticity. Data processing and privacy protection were evaluated in six aspects, including data cleaning and standardization, privacy-preserving mechanisms, and data formatting protocols. As shown in Figure 17, the 53 benchmarks achieved an average score rate of 74.6% (16.4 out of 22) across 11 evaluation criteria, with a mean of 1.4 points per item. Overall, the results indicate general compliance with the proposed standards, although two specific criteria showed suboptimal performance.
Figure 17: Scoring Performance of the 53 Benchmarks in the Phase II: Dataset Construction and Management
@@ B.6 Statistics About Phase II: Dataset Construction and Management
The coverage of each benchmark after calculation is shown in Fig. 18. We scored each benchmark based on the average total coverage rate (21.8%), and the results are presented in Fig 19, 38% (20 out of 53) of the benchmarks lack a clear definition or discussion of dataset diversity. The absence of well-defined diversity goals and corresponding quantitative analyses hinders the ability to assess model robustness across different disease categories or clinical settings, potentially obscuring vulnerabilities in specific subgroups or task scenarios.
@@
Figure 18: Coverage of disease types, medical departments, and overall performance across 53 medical benchmarks.
@@
Prevention of Data Contamination. This dimension assesses whether the benchmark includes mechanisms to identify and mitigate potential data contamination—i.e., the presence of evaluation samples within the training corpus of LLMs. Such contamination can artificially inflate model performance through memorization, thereby compromising the fairness and validity of the evaluation. If present in high-risk tasks (e.g., diagnostic decisions or medication recommendations), contamination may mislead researchers and regulators in their assessment of model safety. Therefore, contamination detection is a prerequisite for ensuring the credibility and scientific rigor of benchmark evaluations. However, as depicted in Figure 20, only 8% (4 out of 53) of the benchmarks implemented both detection and mitigation strategies. Another 49% (26 out of 53) conducted detection without addressing identified contamination, while 43% (23 out of 53) did not mention the issue at all.
@@
B.7 Statistics About Phase III: Technical Implementation and Evaluation Methodology
This is the operational backbone. It consists of 8 items, encompassing the development of accessible, reproducible evaluation scripts and the selection of metrics that go beyond simple accuracy to assess deeper capabilities such as reasoning, robustness, and uncertainty awareness. For the 53 evaluated benchmarks, the average score is 52.4% (8.38 out if 16), with a mean score of 1.05 for each item. Results indicate that this is the second worst performing stage among the 5 stages. As shown in Figure 21, there are 3 items exhibiting critical underperformance, where majority of the benchmarks receive 0 mark.
Figure 21: Scoring Performance of the 53 Benchmarks in the Phase III: Technical Implementation and Evaluation Methodology
@@ B.7 Statistics About Phase III: Technical Implementation and Evaluation Methodology
As shown in Figure 22, a concerning 51% (27 out of 53) of benchmarks do not have any consideration for the models’ reasoning process. While 24% (13 out of 53) of benchmarks mention the importance of evaluating the reasoning process, only 25% (13 out of 53) of benchmarks design concrete assessment for reasoning process with clear methods and metrics.
@@
As depicted in Figure 23, an overwhelming 87% (46 out of 53) of the covered benchmarks do not consider robustness in their evaluation. While 4 benchmarks recognize the importance of robustness evaluation, only 3 benchmarks design clear assessment for robustness.
@@
As illustrated in Figure 24, 73% (39 out of 53) benchmarks have no considerations for uncertainty in their evaluation. Even though 23% (12 out of 53) acknowledge the importance of uncertainty evaluation, only 4% (2 out of 53) incorporate evaluation related to uncertainty in the benchmark, highlighting a significant gap in current benchmarking practices.
@@
B.8 Statistics About Phase IV: Benchmark Validity and Performance Verification
This phase provides scientific validation for the benchmark. It encompasses evidence of content validity (i.e., whether the benchmark comprehensively covers the target domain) and construct validity (i.e., whether it measures the claimed latent capabilities), as well as its ability to distinguish performance differences among models of varying proficiency levels. Among the 53 benchmarks evaluated, the average score rate for this phase was 48.8% (5.9 out of 12), with a mean of 0.5 points per item—representing the lowest performance among the five phases. As demonstrated in Figure 25, of the six evaluation criteria in this phase, three demonstrated notably poor results.
Figure 25: Scoring Performance of the 53 Benchmarks in the Phase IV: Benchmark Validity and Performance Verification
@@ B.8 Statistics About Phase IV: Benchmark Validity and Performance Verification
As presented in Figure 26, among the 53 benchmarks reviewed, 45% (24 benchmarks) did not address the correlation between evaluation scores and clinical performance. A further 53% (28 benchmarks) discussed expected correlations at a theoretical level but lacked supporting experimental evidence. Only 2% (1 benchmark) provided preliminary empirical analysis exploring the correlation, along with a discussion of the results. These findings indicate a significant gap in validating the practical relevance of benchmark scores. Future medical benchmarks should prioritize establishing and verifying correlations between evaluation outcomes and clinical workflows or expert assessments. This is essential to ensure that benchmark scores serve as credible and informative indicators for real-world model deployment.
@@
Internal Consistency. This metric examines whether different components or items within a benchmark consistently measure the same target capability, thereby ensuring internal reliability. High internal consistency reduces the risk that variability in item quality, phrasing, or task context will distort evaluation results. Figure 27 shows that in this study, only 4% (2 out of 53) of the benchmarks reported statistical indicators of internal consistency, with positive results. An additional 24% (13 out of 53) also reported such metrics, but without clear discussion or yielded inconclusive findings. The remaining 72% (38 out of 53) did not conduct internal consistency assessments. The absence of such analysis raises concerns regarding evaluation accuracy and fairness. Inconsistent benchmarks may obscure true model strengths or weaknesses across specific skill dimensions, impeding targeted model refinement. In clinical applications, this may further compromise risk assessment and regulatory oversight. Therefore, future benchmarks should prioritize the evaluation and reporting of internal consistency to enhance reliability.
@@
It can be seen from Figure 28 that, in practice, 59% (31 out of 53) of the benchmarks reported only single-run results. Although 32% (17 out of 53) provided mean and standard deviation across multiple runs, they did not employ formal statistical tests to compare model performance. These findings highlight a lack of rigorous statistical standards in current benchmark design. Incorporating structured multi-run strategies and statistical significance testing is critical to enhance the reliability and validity of performance evaluations in medical AI.
@@
B.9 Statistics About Phase V: Documentation, Openness, and Governance
Documentation, Openness, and Governance represent community-facing components that are essential for ensuring the long-term value and impact of a benchmark. This phase includes the provision of clear documentation, adherence to open-source principles, and the establishment of governance mechanisms for maintenance, version control, and community feedback. Among the 53 benchmarks evaluated, as depicted in Figure 29, the average score rate for this phase was 65.2% (11.7 out of 18), with a mean score of 1.0 per item. The scoring distribution in this phase was relatively balanced, with only one criterion showing slightly weaker performance.
Figure 29: Scoring Performance of the 53 Benchmarks in the Phase V: Documentation, Openness, and Governance
Discussion of Limitations and Risks. This criterion evaluates whether the benchmark documentation openly and explicitly discusses its inherent limitations, potential societal risks, or misuse scenarios. No benchmark is flawless. Proactively disclosing such limitations and risks reflects scientific rigor and a strong sense of responsibility on the part of the developers. It also helps prevent overreliance on benchmark results or inappropriate applications. Figure 30 shows that Only 7% (4 out of 53) of the benchmarks provided thorough and candid discussions of limitations and associated risks. A majority—74% (39 out of 53)—briefly mentioned limitations but lacked in-depth analysis. The remaining 19% (10 out of 53) did not discuss any form of limitation or risk.

### v2

@@ Appendix F Quantitative Results
This appendix provides a summary of the quantitative analysis conducted on the 56 medical LLM benchmarks.
F.1 Publication Year and Task Focus of Benchmarks
@@
Annual Trend. As shown in Figure 4, the publication years of the 56 benchmarks we collected. A clear turning point can be observed in 2024, when the volume of medical benchmark publications significantly increased. Prior to 2024, only 17 benchmarks were released. The brief surge in 2019, driven by the rise of deep learning and Transformer models, catalyzed short-term growth in medical AI and a peak in benchmark publications. However, no substantial upward trend was evident before 2024.
Figure 4: Publication Year Distribution of the 56 Medical Benchmarks
In contrast, with the emergence of large models such as ChatGPT and Med-PaLM in 2024, there was an explosive growth in medical NLP and multimodal benchmarks. A total of 27 medical benchmarks were published in 2024, surpassing the cumulative number of the previous six years. As of July 10, 2025, 12 new benchmarks were already released in the first half of 2025, showing a steady growth trend.
@@
In summary, as illustrated in Figure 6, more than half (64%) of the existing medical benchmarks focus on assessing models’ text comprehension, NLP, and VQA abilities. This indicates that most medical benchmarks historically centered around structured or unstructured medical text understanding. However, in recent years, there has been an increasing trend towards medical question answering and clinical NLP, with the highest number of such benchmarks being published in the first half of 2025.
Figure 6: Task Type Distribution of the 56 Medical Benchmarks
F.2 Natural Language
Figure 7 presents the distribution of natural language usage across the 56 collected benchmarks. It is evident that English dominates, being used in 45 of the benchmarks (approximately 80%), establishing itself as the primary language for evaluating medical LLMs. Chinese ranks second, appearing in 17 benchmarks (approximately 30%), indicating a notable level of development in Chinese medical AI. In contrast, other languages such as Arabic, Russian, Portuguese, Japanese, and Hindi—despite being spoken by large global populations—remain underrepresented in publicly available benchmarks, with each language appearing in only 2 to 6 benchmarks on average. This highlights that multilingual medical AI remains in its early stages.
@@ F.2 Natural Language
F.3 Medical Diversity
Distribution of Diseases. Figure 8 illustrates the distribution of disease categories represented in the 56 medical benchmarks, categorized according to the first 23 chapters of the ICD-11 classification system (with Certain conditions originating in the perinatal period and Developmental anomalies excluded due to the absence of relevant benchmarks). The analysis reveals that current research on medical LLMs is heavily concentrated in disease areas such as Diseases of the musculoskeletal system or connective tissue (16 benchmarks), Diseases of the skin (15), Diseases of the respiratory system (14), and Neoplasms (13). These categories typically have abundant publicly available imaging data (e.g., X-rays, CT scans, pathology slides), which facilitates the development of standardized evaluation protocols based on both image and text modalities.
@@ F.3 Medical Diversity
F.4 Benchmark Performance Heatmap Example
@@
The heatmap in Figure 10 visualizes the performance of 14 out of 56 of the medical LLM benchmarks evaluated evaluated against the five-phase MedCheck framework. Each row represents a benchmark, and columns correspond to the framework’s phases. The color intensity reflects the score rate (%), highlighting the comparative strengths and weaknesses of each benchmark and revealing systemic trends.
F.5 Statistics About Phase I: Design and Conceptualization
@@
As shown in Figure 11, the average normalized score across benchmarks is 74.5% (17.9 out of 24), with a mean of 1.5 points per criterion, indicating general compliance with baseline expectations. However, performance on several key items remains suboptimal and warrants further attention.
Figure 11: Scoring Performance of the 56 Benchmarks in the Design and Conceptualization Stage
Clearly specified application scenarios. To ensure practical relevance, benchmarks should clearly define their intended clinical or biomedical research scenarios. Such grounding facilitates real-world applicability and aligns evaluation with deployment needs. As illustrated in Figure 12, 68% of benchmarks specify concrete application scenarios, articulating clinical value and expected utility. In contrast, 18% provide only general references to decision-support contexts, and 14% fail to specify any scenario, reducing interpretability and downstream usability.
@@
Involvement of domain experts. Expert input is a critical component in medical AI development. It ensures the appropriateness of task design, data annotation quality, and answer validity. However, as shown in Figure 13, only 23% of benchmarks explicitly report domain expert involvement, often without detailing the roles or qualifications. 16% do not mention expert participation, raising concerns about the benchmarks’ clinical validity. Future benchmark design should adopt standardized expert reporting practices to enhance credibility, transparency, and methodological rigor.
@@
Alignment with medical standards. Adherence to international medical standards (e.g., ICD-11, SNOMED CT, LOINC) is essential for achieving semantic interoperability, annotation consistency, and cross-system comparability. As Figure 14 indicates, 50% of benchmarks do not clarify whether their data conforms to such standards. This omission undermines the benchmark’s reusability and weakens the comparability of evaluation outcomes. Without standard-based structure and terminology, label ambiguity and inconsistent categorization may arise, which compromises the reliability of model performance evaluation.
@@
Multidimensional evaluation design. In medical AI, a model’s performance cannot be solely measured by accuracy. Outputs that are correct but unsafe, incomplete, or biased may have detrimental consequences. Multidimensional evaluation frameworks are required to assess whether models’ output is not only correct but also reliable. As presented in Figure 15, 39% of benchmarks use only a single metric, typically accuracy, and 20% mention other metrics without detailed implementation. Metrics such as uncertainty estimation, answer refusal, and harmful output detection are absent from most benchmarks, despite their relevance in high-stakes medical decision-making.
@@
Consideration of safety and fairness. Robust evaluation frameworks must account for safety and fairness, especially in sensitive domains like healthcare. As shown in Figure 16, only 14% of benchmarks include explicit safety or fairness evaluations, while 36% acknowledge their importance without implementation. The remaining 50% omit these aspects entirely. Neglecting safety assessments increases the risk of misleading or overconfident outputs, which may harm clinical decisions. The lack of fairness evaluation may propagate systematic bias across gender, age, race, or disease distribution, exacerbating healthcare disparities. To build trustworthy and ethical medical AI, safety and fairness must be treated as core evaluation dimensions.
@@
F.6 Statistics About Phase II: Dataset Construction and Management
This phase constitutes the empirical foundation of the benchmark, with an emphasis on constructing datasets that are realistic, diverse, representative, and ethically sourced. Key elements include stringent procedures for data quality assurance, privacy protection, and contamination prevention. The dataset sources and quality were assessed across five dimensions: transparency and traceability of data provenance, reliability of data sources, and data authenticity. Data processing and privacy protection were evaluated in six aspects, including data cleaning and standardization, privacy-preserving mechanisms, and data formatting protocols. As shown in Figure 17, the 56 benchmarks achieved an average score rate of 74.7 % (16.4 out of 22) across 11 evaluation criteria, with a mean of 1.4 points per item. Overall, the results indicate general compliance with the proposed standards, although two specific criteria showed suboptimal performance.
Figure 17: Scoring Performance of the 56 Benchmarks in the Phase II: Dataset Construction and Management
@@ F.6 Statistics About Phase II: Dataset Construction and Management
The coverage of each benchmark after calculation is shown in Fig. 18. We scored each benchmark based on the average total coverage rate (21.8%), and the results are presented in Fig 19, 39% (21 out of 56) of the benchmarks lack a clear definition or discussion of dataset diversity. The absence of well-defined diversity goals and corresponding quantitative analyses hinders the ability to assess model robustness across different disease categories or clinical settings, potentially obscuring vulnerabilities in specific subgroups or task scenarios.
@@
Figure 18: Coverage of disease types, medical departments, and overall performance across 56 medical benchmarks.
@@
Prevention of Data Contamination. This dimension assesses whether the benchmark includes mechanisms to identify and mitigate potential data contamination—i.e., the presence of evaluation samples within the training corpus of LLMs. Such contamination can artificially inflate model performance through memorization, thereby compromising the fairness and validity of the evaluation. If present in high-risk tasks (e.g., diagnostic decisions or medication recommendations), contamination may mislead researchers and regulators in their assessment of model safety. Therefore, contamination detection is a prerequisite for ensuring the credibility and scientific rigor of benchmark evaluations. However, as depicted in Figure 20, only 7% of the benchmarks implemented both detection and mitigation strategies. Another 48% conducted detection without addressing identified contamination, while 45% did not mention the issue at all.
@@
F.7 Statistics About Phase III: Technical Implementation and Evaluation Methodology
This is the operational backbone. It consists of 8 items, encompassing the development of accessible, reproducible evaluation scripts and the selection of metrics that go beyond simple accuracy to assess deeper capabilities such as reasoning, robustness, and uncertainty awareness. For the 56 evaluated benchmarks, the average score is 52.1% (8.33 out of 16), with a mean score of 1.04 for each item. Results indicate that this is the second worst performing stage among the 5 stages. As shown in Figure 21, there are 3 items exhibiting critical underperformance, where majority of the benchmarks receive 0 mark.
Figure 21: Scoring Performance of the 56 Benchmarks in the Phase III: Technical Implementation and Evaluation Methodology
@@ F.7 Statistics About Phase III: Technical Implementation and Evaluation Methodology
As shown in Figure 22, a concerning 55% of benchmarks do not have any consideration for the models’ reasoning process. While 25% of benchmarks mention the importance of evaluating the reasoning process, only 23% of benchmarks design concrete assessment for reasoning process with clear methods and metrics.
@@
As depicted in Figure 23, an overwhelming 88% of the covered benchmarks do not consider robustness in their evaluation. While 4 benchmarks recognize the importance of robustness evaluation, only 3 benchmarks design clear assessment for robustness.
@@
As illustrated in Figure 24, 75% benchmarks have no considerations for uncertainty in their evaluation. Even though 21% acknowledge the importance of uncertainty evaluation, only 4% incorporate evaluation related to uncertainty in the benchmark, highlighting a significant gap in current benchmarking practices.
@@
F.8 Statistics About Phase IV: Benchmark Validity and Performance Verification
This phase provides scientific validation for the benchmark. It encompasses evidence of content validity (i.e., whether the benchmark comprehensively covers the target domain) and construct validity (i.e., whether it measures the claimed latent capabilities), as well as its ability to distinguish performance differences among models of varying proficiency levels. Among the 56 benchmarks evaluated, the average score rate for this phase was 49.1% (5.90 out of 12), with a mean of 0.98 points per item—representing the lowest performance among the five phases. As demonstrated in Figure 25, of the six evaluation criteria in this phase, three demonstrated notably poor results.
Figure 25: Scoring Performance of the 56 Benchmarks in the Phase IV: Benchmark Validity and Performance Verification
@@ F.8 Statistics About Phase IV: Benchmark Validity and Performance Verification
As presented in Figure 26, among the 56 benchmarks reviewed, 45% did not address the correlation between evaluation scores and clinical performance. A further 53% (28 benchmarks) discussed expected correlations at a theoretical level but lacked supporting experimental evidence. Only 2% (1 benchmark) provided preliminary empirical analysis exploring the correlation, along with a discussion of the results. These findings indicate a significant gap in validating the practical relevance of benchmark scores. Future medical benchmarks should prioritize establishing and verifying correlations between evaluation outcomes and clinical workflows or expert assessments. This is essential to ensure that benchmark scores serve as credible and informative indicators for real-world model deployment.
@@
Internal Consistency. This metric examines whether different components or items within a benchmark consistently measure the same target capability, thereby ensuring internal reliability. High internal consistency reduces the risk that variability in item quality, phrasing, or task context will distort evaluation results. Figure 27 shows that in this study, only 4% of the benchmarks reported statistical indicators of internal consistency, with positive results. An additional 23% also reported such metrics, but without clear discussion or yielded inconclusive findings. The remaining 73% did not conduct internal consistency assessments. The absence of such analysis raises concerns regarding evaluation accuracy and fairness. Inconsistent benchmarks may obscure true model strengths or weaknesses across specific skill dimensions, impeding targeted model refinement. In clinical applications, this may further compromise risk assessment and regulatory oversight. Therefore, future benchmarks should prioritize the evaluation and reporting of internal consistency to enhance reliability.
@@
It can be seen from Figure 28 that, in practice, 57% of the benchmarks reported only single-run results. Although 30% provided mean and standard deviation across multiple runs, they did not employ formal statistical tests to compare model performance. These findings highlight a lack of rigorous statistical standards in current benchmark design. Incorporating structured multi-run strategies and statistical significance testing is critical to enhance the reliability and validity of performance evaluations in medical AI.
@@
F.9 Statistics About Phase V: Documentation, Openness, and Governance
Documentation, Openness, and Governance represent community-facing components that are essential for ensuring the long-term value and impact of a benchmark. This phase includes the provision of clear documentation, adherence to open-source principles, and the establishment of governance mechanisms for maintenance, version control, and community feedback. Among the 56 benchmarks evaluated, as depicted in Figure 29, the average score rate for this phase was 66.2% (11.9 out of 18), with a mean score of 1.3 per item. The scoring distribution in this phase was relatively balanced, with only one criterion showing slightly weaker performance.
Figure 29: Scoring Performance of the 56 Benchmarks in the Phase V: Documentation, Openness, and Governance
Discussion of Limitations and Risks. This criterion evaluates whether the benchmark documentation openly and explicitly discusses its inherent limitations, potential societal risks, or misuse scenarios. No benchmark is flawless. Proactively disclosing such limitations and risks reflects scientific rigor and a strong sense of responsibility on the part of the developers. It also helps prevent overreliance on benchmark results or inappropriate applications. Figure 30 shows that Only 11% of the benchmarks provided thorough and candid discussions of limitations and associated risks. A majority—71%—briefly mentioned limitations but lacked in-depth analysis. The remaining 18% (10 out of 53) did not discuss any form of limitation or risk.